OpenStudy (anonymous):

how would you solve c^2+9c+18=0 because i think its non-factorable?

OpenStudy (campbell_st):

there are 2 methods 1 completing the square which requires rewriting the equation \[c^2 + 9c = -18\] then finding the value of (9/2)^2 and adding it to both sides \[c^2 + 9c + \frac{81}{4} = -18 + \frac{81}{4}\] left hand side is a perfect square and the RHS is just a number 2 using the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

thanks dude i forgot about the general quadratic formula its easier for me to use that thanks